# Causal Graphical Models for Vision-Language Compositional Understanding

### v1

@@ Abstract
Recent work has empirically shown that Vision-Language Models (VLMs) struggle to fully understand the compositional properties of the human language, usually modeling an image caption as a “bag of words”. As a result, they perform poorly on compositional tasks, which require a deeper understanding of the different entities of a sentence (subject, verb, etc.) jointly with their mutual relationships in order to be solved. In this paper, we model the dependency relations among textual and visual tokens using a Causal Graphical Model (CGM), built using a dependency parser, and we train a decoder conditioned by the VLM visual encoder. Differently from standard autoregressive or parallel predictions, our decoder’s generative process is partially-ordered following the CGM structure. This structure encourages the decoder to learn only the main causal dependencies in a sentence discarding spurious correlations. Using extensive experiments on five compositional benchmarks, we show that our method significantly outperforms all the state-of-the-art compositional approaches by a large margin, and it also improves over methods trained using much larger datasets.  [^1] Our code and our models are available at: https://github.com/aimagelab/COGT
@@ 1 Introduction
Vision-Language Models (VLMs) have shown impressive results in different tasks such as, for instance, zero-shot classification, image-text retrieval, vision-question answering, image-captioning, and many others (Radford et al., 2021; Li et al., 2023b; Singh et al., 2021; Liu et al., 2023). However, despite this success, most VLMs still struggle in understanding the compositional nature of the human language. For instance, Yuksekgonul et al. (2023) empirically showed that common VLMs usually do not consider the order and the syntactic/semantic relations of words in a sentence, which is treated as a bag of words, where “the horse is eating the grass” and “the grass is eating the horse” can easily be confused. Jointly with Yuksekgonul et al. (2023), many other authors have recently proposed different compositional benchmarks which confirm the poor performance of common VLMs when tested against compositional tasks (Hsieh et al., 2024; Zhao et al., 2022; Burapacheep et al., 2024). One of the probable reasons of this bag-of-words behavior is the contrastive loss used in CLIP (Radford et al., 2021) (and in other VLMs), which compares a single vector representing the textual encoder’s output with a single vector representing the visual encoder’s output, sacrificing textual and visual details (Yuksekgonul et al., 2023; Kamath et al., 2023; Basu et al., 2024). Another reason is the low quality of the captions used for VLM pre-training, which are usually noisy or do not describe the details of the image and the interactions among its objects (Doveh et al., 2023a).
Most of the compositional methods that have been recently proposed to alleviate this problem focus on creating annotations with a richer compositional structure, used to fine-tune a VLM (Yuksekgonul et al., 2023; Doveh et al., 2023a; Cascante-Bonilla et al., 2023). For instance, NegCLIP (Yuksekgonul et al., 2023) creates hard negatives, in which the original caption is modified swapping the positions of some words, and these hard negatives are used jointly with common negatives to fine-tune CLIP using the standard contrastive loss. However, the automatic creation of hard negatives is itself noisy, leading to captions which often do not have a correct syntactic/semantic meaning (this problem is inherited by some compositional benchmarks, see Sec. 4). In (Tschannen et al., 2023), a VLM is pre-trained from scratch using a captioning strategy and a huge private dataset. Specifically, the authors propose both Cap, where the pre-training strategy is a standard AutoRegressive (AR) next-token prediction, and CapPa, where the AR training is mixed with a parallel training, in which all the textual tokens are simultaneously predicted. Tschannen et al. (2023) show that both Cap and CapPa achieve excellent results on compositional tasks, and argue that a generative training encourages the VLM to focus on fine-grained descriptions of the visual content.
@@ 2 Related work
Syntactic Trees. In Dependency Grammars, dependency relations are syntactic and semantic connections between words in a sentence, where one word (called “head”) governs or determines the grammatical behavior of the “dependent” word (Nivre, 2005). Given a sentence, these dependencies are organized in a Dependency Tree (DT), which can be automatically extracted using a parser (Honnibal et al., 2020; Zhang et al., 2020; Dozat & Manning, 2016). In (Yang & Wan, 2022) a Language Model (LM) is trained to predict whether the future tokens in the sequence are head or dependent of a previously observed token, and this prediction replaces the standard Maximum Likelihood Estimation objective. Similarly, in (Deguchi et al., 2019) the dependency relations extracted by an external parser are learned by the LM and used to modulate the Transformer (Vaswani et al., 2017) attention maps. A DT can also be used to compute a syntactic distance between words, which in turn can be used, e.g., as an additional loss (Du et al., 2020) or to modulate the attention maps (Hou et al., 2022).
In the vision-language domain, DTs are used in (Song et al., 2022) to convert a textual question into a template for CLIP fine-tuning, and in (Li et al., 2024b) to replace and swap words in a sentence and fine-tune a VLM using the correction of the modified sequence as a pretext task. Finally, constituency parsers group words that belong to a specific grammatical category in a sub-phrase. Constituency trees are used, e.g., in (Yellinek et al., 2023) to generate sub-phrase specific hard negative captions or in (Zhang et al., 2022) to extended the contrastive loss by maximizing the similarity of words in the same sub-phrase. Differently from previous work, we use a DT to extract syntactic and semantic dependencies between the words of a sentence, and we interpret these dependencies as causal relations that guide the construction of our CGM.
@@ 4 Experiments
Mixed 84.83 69.12 76.98 99.01 84.17 78.39 87.19 85.89 76.96 91.44 84.76 41.33 45.21       67.10${}_{\textbf{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+9.67}}}$
Sequential-AR 84.86 77.87 81.37 98.96 83.62 81.50 88.02 86.82 75.45 91.11 84.45 46.33 46.24       69.28${}_{\textbf{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+2.18}}}$
COGT 87.56 90.26 88.91 98.26 87.10 83.14 89.50 86.07 78.91 89.37 84.78 61.33 51.48       75.20${}_{\textbf{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}+5.92}}}$
@@ Acknowledgments
This work was supported by the PNRR project Italian Strengthening of Esfri RI Resilience (ITSERR) funded by the European Union – NextGenerationEU (CUP B53C22001770006), and by the EU Horizon project ELIAS (No. 101120237).
Appendix A Causal Grafical Models

### v2

@@ Abstract
Recent work has empirically shown that Vision-Language Models (VLMs) struggle to fully understand the compositional properties of the human language, usually modeling an image caption as a “bag of words”. As a result, they perform poorly on compositional tasks, which require a deeper understanding of the different entities of a sentence (subject, verb, etc.) jointly with their mutual relationships in order to be solved. In this paper, we model the dependency relations among textual and visual tokens using a Causal Graphical Model (CGM), built using a dependency parser, and we train a decoder conditioned by the VLM visual encoder. Differently from standard autoregressive or parallel predictions, our decoder’s generative process is partially-ordered following the CGM structure. This structure encourages the decoder to learn only the main causal dependencies in a sentence discarding spurious correlations. Using extensive experiments on five compositional benchmarks, we show that our method significantly outperforms all the state-of-the-art compositional approaches by a large margin, and it also improves over methods trained using much larger datasets. Our model weights and code are publicly available. [^1] https://aimagelab.github.io/COGT
@@ 1 Introduction
Vision-Language Models (VLMs) have shown impressive results in different tasks such as, for instance, zero-shot classification, image-text retrieval, vision-question answering, image-captioning, and many others (Radford et al., 2021; Li et al., 2023b; Singh et al., 2021; Liu et al., 2023). However, despite this success, most VLMs still struggle in understanding the compositional nature of the human language. For instance, Yuksekgonul et al. (2023) empirically showed that common VLMs usually do not consider the order and the syntactic/semantic relations of words in a sentence, which is treated as a bag of words, where “the horse is eating the grass” and “the grass is eating the horse” can easily be confused. Jointly with Yuksekgonul et al. (2023), many other authors have recently proposed different compositional benchmarks which confirm the poor performance of common VLMs when tested against compositional tasks (Hsieh et al., 2024; Zhao et al., 2022; Burapacheep et al., 2024; Peng et al., 2024). One of the probable reasons of this bag-of-words behavior is the contrastive loss used in CLIP (Radford et al., 2021) (and in other VLMs), which compares a single vector representing the textual encoder’s output with a single vector representing the visual encoder’s output, sacrificing textual and visual details (Yuksekgonul et al., 2023; Kamath et al., 2023; Basu et al., 2024). Another reason is the low quality of the captions used for VLM pre-training, which are usually noisy or do not describe the details of the image and the interactions among its objects (Doveh et al., 2023a).
Most of the compositional methods that have been recently proposed to alleviate this problem focus on creating annotations with a richer compositional structure, used to fine-tune a VLM (Yuksekgonul et al., 2023; Doveh et al., 2023a; Cascante-Bonilla et al., 2023). For instance, NegCLIP (Yuksekgonul et al., 2023) creates hard negatives, in which the original caption is modified swapping the positions of some words, and these hard negatives are used jointly with common negatives to fine-tune CLIP using the standard contrastive loss. However, the automatic creation of hard negatives is itself noisy, leading to captions which often do not have a correct syntactic/semantic meaning (this problem is inherited by some compositional benchmarks, see Sec. 4). In (Tschannen et al., 2023), a VLM is pre-trained from scratch using a captioning strategy and a huge private dataset. Specifically, the authors propose both Cap, where the pre-training strategy is a standard AutoRegressive (AR) next-token prediction, and CapPa, where the AR training is mixed with a parallel training (Bolelli et al., 2018), in which all the textual tokens are simultaneously predicted. Tschannen et al. (2023) show that both Cap and CapPa achieve excellent results on compositional tasks, and argue that a generative training encourages the VLM to focus on fine-grained descriptions of the visual content.
@@ 2 Related work
Syntactic Trees. In Dependency Grammars, dependency relations are syntactic and semantic connections between words in a sentence, where one word (called “head”) governs or determines the grammatical behavior of the “dependent” word (Nivre, 2005). Given a sentence, these dependencies are organized in a Dependency Tree (DT), which can be automatically extracted using a parser (Honnibal et al., 2020; Zhang et al., 2020; Dozat & Manning, 2016). In (Yang & Wan, 2022) a Language Model (LM) is trained to predict whether the future tokens in the sequence are head or dependent of a previously observed token, and this prediction replaces the standard Maximum Likelihood Estimation objective. Similarly, in (Deguchi et al., 2019) the dependency relations extracted by an external parser are learned by the LM and used to modulate the Transformer (Vaswani et al., 2017) attention maps. A DT can also be used to compute a syntactic distance between words, which in turn can be used, e.g., as an additional loss (Du et al., 2020) or to modulate the attention maps (Hou et al., 2022). In the vision-language domain, DTs are used in (Song et al., 2022) to convert a textual question into a template for CLIP fine-tuning, and in (Li et al., 2024b) to replace and swap words in a sentence and fine-tune a VLM using the correction of the modified sequence as a pretext task. Finally, constituency parsers group words that belong to a specific grammatical category in a sub-phrase. Constituency trees are used, e.g., in (Yellinek et al., 2023) to generate sub-phrase specific hard negative captions or in (Zhang et al., 2022) to extended the contrastive loss by maximizing the similarity of words in the same sub-phrase. Differently from previous work, we use a DT to extract syntactic and semantic dependencies between the words of a sentence, and we interpret these dependencies as causal relations that guide the construction of our CGM.
@@ 4 Experiments
Mixed 84.83 69.12 76.98 99.01 84.17 78.39 87.19 85.89 76.96 91.44 84.76 41.33 45.21       67.10${}_{\textbf{{\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}+9.67}}}$
Sequential-AR 84.86 77.87 81.37 98.96 83.62 81.50 88.02 86.82 75.45 91.11 84.45 46.33 46.24       69.28${}_{\textbf{{\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}+2.18}}}$
COGT 87.56 90.26 88.91 98.26 87.10 83.14 89.50 86.07 78.91 89.37 84.78 61.33 51.48       75.20${}_{\textbf{{\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}+5.92}}}$
@@ Acknowledgments
This work was supported by the PNRR project “Italian Strengthening of Esfri RI Resilience (ITSERR)”, funded by the European Union – NextGenerationEU (CUP B53C22001770006), by the EU Horizon project ELIAS (No. 101120237), and by the PNRR project “A Static and Dynamic Database for Historic Urban Contexts Accessibility”, funded by the European Union – NextGenerationEU (CUP E53D23010500001).
Appendix A Causal Graphical Models